### v1

@@ 1 Introduction
Since the top quark is by far the heaviest quark in the SM, it also has the largest impact on the cross section, hence Higgs couplings to lighter quarks are often neglected. Moreover, the top quark’s large mass enables the use of the heavy-top limit (HTL) approximation, where the top quark is treated as infinitely massive. This method is valid up to power corrections of order $m_{H}^{2}/4m_{t}^{2}$, and significantly streamlines computations by reducing the loop order by one and removing the top-quark mass as a scale. These simplifications have allowed for calculations at next-to-leading Dawson:1990zj ; Djouadi:1991tka , next-to-next-to-leading Ravindran:2002dc ; Harlander:2002wh ; Anastasiou:2002yz , and next-to-next-to-next-to-leading Anastasiou:2015vya ; Mistlberger:2018etf ; Anastasiou:2016cez order in $\alpha_{s}$ (NLO, NNLO, $\text{N}^{3}$), yielding state-of-the-art predictions with a theoretical uncertainty of approximately $6\%$. Studies have shown that the gluon-gluon-fusion channel receives particularly large perturbative corrections. In fact, the cross section is more than doubled going from LO to NLO, and NNLO corrections remain substantial, contributing roughly 20% to the total cross section. Because of the large size of the corrections, accounting for the previously omitted finite top-quark mass is strongly advisable also at higher perturbative orders. In Higgs effective field theory (HEFT), finite-top-quark-mass effects are included by rescaling the HTL cross section to recover the LO result with finite top-quark mass:
@@
Finite-quark-mass effects can also be leveraged to determine the Yukawa couplings to the respective quarks. The sensitivity of the measurements can be significantly enhanced by additionally considering differential cross sections, e.g. the Higgs-$p_{T}$ Bishara:2016jga ; Bonner:2016sdg or the Higgs-rapidity Soreq:2016rae distribution. This is particularly useful for lighter quarks like the charm quark. CMS studies CMS:2018gwt ; CMS:2023gjz have found an upper bound for the charm Yukawa coupling of about five times its SM value, by fitting NLO predictions to the measured distributions. Meanwhile, NNLO predictions for $p_{T}$ distributions with finite top and bottom mass have become available Grazzini:2013mca ; Melnikov:2016emg ; Lindert:2017pky ; Caola:2018zye ; Bonciani:2022jmb and could potentially yield improved bounds. Rapidity distributions on the other hand are even more challenging, as they require a contribution from the three-loop virtual corrections absent in the $p_{T}$ distribution. Our study aims to bridge this gap, potentially enhancing the sensitivity for future analyses of this kind.
@@ 2 Computational Details
with $\hat{s},\hat{t}$ and $\hat{u}$ the usual partonic Mandelstam variables, remains challenging. In order to minimise the computational effort, a particular basis of master integrals involving dimension shifts for the most complicated integrals is chosen. Furthermore, we use a fixed ratio of the quark and Higgs mass square, $m_{q}^{2}/m_{H}^{2}$, to eliminate one of the scales in favour of a rational number, which further simplifies the rational-function coefficients in front of the master integrals. In fact, we compute the master integrals and the complete amplitudes for five different bottom-quark masses and four different top-quark masses, listed in tab. 1. After the reduction to master integrals, we derive a set of coupled differential equations. The integrals for the boundary condition at $x\rightarrow\infty$ can be solved with the help of the large-mass expansion. We then transport the solution to physical values of $x$, by solving the differential equations numerically using the Bulirsch-Stoer algorithm (see fig. 1).
@@
This is done multiple times for many values of $z$ and a fixed value of $\lambda$ close to the symmetry axis of $\lambda=1/2$. Subsequently, we solve the differential equations in $\lambda$ to map out the entire physical plane, marked red in fig. 1, such that we obtain the amplitude as a dense two dimensional grid. During the Monte Carlo integration we interpolate between the grid points using cubic splines. The accuracy of the splines was verified by removing a mass grid and recalculating the resulting cross sections. The discrepancies between the recalculated cross sections and the original ones were below the Monte Carlo uncertainties, confirming the robustness of our interpolation method. We note that we only calculated the interference of all bottom-quark amplitudes with the OS-renormalised, as well as with the $\overline{\mathrm{MS}}$-renormalised top-quark amplitudes at the same renormalisation scale, to construct grids for cubic-spline interpolation.
@@
To further improve the quality of the spline fits, we subtract all infrared divergences from the amplitudes and add them back only after performing the interpolation. The subtraction of infrared divergences can be simplified by first subtracting a rescaled cross section of the HTL,
@@
$\begin{split}&\braket{M^{(0)}_{gg\rightarrow Hg}}{M^{(1)}_{gg\rightarrow Hg}}|_{\text{regulated}}\equiv\braket{M^{(0)}_{gg\rightarrow Hg}}{M^{(1)}_{gg\rightarrow Hg}}\\ &\quad-\left[r\braket{M^{(0)}_{gg\rightarrow Hg,\text{HTL}}}{M^{(1)}_{gg\rightarrow Hg,\text{HTL}}}+\frac{8\pi\alpha_{s}}{\hat{t}}\braket{P_{gg}^{(0)}\!\left(\frac{\hat{s}}{\hat{s}+\hat{u}}\right)}\braket{M^{(0)}_{gg\rightarrow H}}{M^{(1)}_{gg\rightarrow H}-M^{(1)}_{gg\rightarrow H,\text{HTL}}}\right]\\ &\braket{P_{gg}^{(0)}\!\left(z\right)}=2C_{A}\left(\frac{z}{1-z}+\frac{1-z}{z}+z(1-z)\right).\end{split}$ (4)
The rescaling factor is chosen in such a way as to match the LO cross section with finite quark mass
$r=\frac{\sigma^{\text{LO}}_{i}}{\sigma^{\text{LO}}_{\text{HTL}}},$ (5)
where $\sigma_{i}$ refers to the cross section of the interference contribution or the cross section with finite top-quark mass. This way, the first subtraction term in eq. (4) already captures most of the infrared singularities. The only infrared divergence which is not yet removed stems from a soft or collinear splitting of the two-loop form factors. After subtracting these left-over divergences with the corresponding splitting functions, the regulated squared matrix element is free of infrared singularities apart from logarithmic ones. Note that the splitting function also removes the soft divergences. The regulated squared amplitude for $qg\longrightarrow qH$ is defined analogously with the above splitting function replaced by
$-\braket{P_{qq}^{(0)}(z)}=-T_{F}\left(1-2z(1-z)\right).$ (6)
The amplitude for $q\bar{q}\longrightarrow gH$, on the other hand, has no additional singularities and no subtraction of splitting functions is needed.
@@ 2.1 $\overline{\text{MS}}$ scheme
$Z_{m}^{\overline{\text{MS}}}\overline{m}=Z_{m}^{\text{OS}}m^{\text{OS}},$ (7)
@@
$m^{\text{OS}}=\overline{m}\left(1+c_{1}\frac{\alpha_{s}}{\pi}+c_{2}\left(\frac{\alpha_{s}}{\pi}\right)^{2}+\mathcal{O}(\alpha^{3})\right).$ (8)
@@
$\begin{split}c_{1}&=C_{F}\left(1-\frac{3}{4}\ln\!\frac{\overline{m}^{2}}{\mu^{2}}\right),\\ c_{2}&=-\frac{1}{384}C_{F}\bigg{(}C_{A}\left(-132\log^{2}\!\frac{\overline{m}^{2}}{\mu^{2}}+740\log\!\frac{\overline{m}^{2}}{\mu^{2}}+144\zeta(3)-1111+\pi^{2}(32-96\log(2))\right)\\ &\hskip 28.45274pt+3C_{F}\bigg{(}-36\log^{2}\!\frac{\overline{m}^{2}}{\mu^{2}}-36\log\!\frac{\overline{m}^{2}}{\mu^{2}}-96\zeta(3)+71+8\pi^{2}(8\log(2)-5)\bigg{)}\\ &\hskip 28.45274pt+4n_{l}T_{F}\left(12\log^{2}\!\frac{\overline{m}^{2}}{\mu^{2}}-52\log\!\frac{\overline{m}^{2}}{\mu^{2}}+8\pi^{2}+71\right)\bigg{)}\\ &\quad-\sum_{i}\frac{1}{96}C_{F}T_{F}\bigg{(}12\log^{2}\!\frac{\overline{m}^{2}}{\mu^{2}}-24\log\!\frac{\overline{m}^{2}}{\mu^{2}}\log\!\frac{\overline{m}_{i}^{2}}{\mu^{2}}-52\log\!\frac{\overline{m}^{2}}{\mu^{2}}+32\log\!\frac{\overline{m}_{i}^{2}}{\mu^{2}}\\ &\hskip 28.45274pt+48(x_{i}-1)^{2}\left(x_{i}^{2}+x_{i}+1\right)\text{Li}_{2}(1-x_{i})-48\left(x_{i}^{4}+x_{i}^{3}+x_{i}+1\right)\left(\text{Li}_{2}(-x_{i})+\log\!(x_{i})\log\!(x_{i}+1)\right)\\ &\hskip 28.45274pt+48x_{i}^{4}\log^{2}(x_{i})+48x_{i}^{2}\log(x_{i})-16\pi^{2}(x_{i}^{3}+x_{i})+72x_{i}^{2}+71\bigg{)},\end{split}$ (10)
where we defined $x_{i}=\frac{\overline{m}_{i}}{\overline{m}}$. The sum runs over all heavy-quark flavours which are decoupled from the running of $\alpha_{s}$. This includes $\overline{m}$ itself in case the corresponding quark is decoupled. Subsequently, the amplitudes can simply be expanded to find the relation between the two schemes:
$\begin{split}&\mathcal{M}^{\overline{\text{MS}}}=\mathcal{M}^{\text{OS}}+\delta\mathcal{M},\\ &\delta\mathcal{M}^{(1)}=\overline{m}c_{1}\frac{\alpha_{s}}{\pi}\frac{\mathrm{d}\mathcal{M}^{\text{OS},(0)}}{\mathrm{d}m}\bigg{|}_{m=\overline{m}},\\ &\delta\mathcal{M}^{(2)}=\overline{m}\left[c_{1}\frac{\alpha_{s}}{\pi}\frac{\mathrm{d}\mathcal{M}^{\text{OS},(1)}}{\mathrm{d}m}\bigg{|}_{m=\overline{m}}+c_{2}\left(\frac{\alpha_{s}}{\pi}\right)^{2}\frac{\mathrm{d}\mathcal{M}^{\text{OS},(0)}}{\mathrm{d}m}\bigg{|}_{m=\overline{m}}\right]+\frac{1}{2}\left(\overline{m}c_{1}\frac{\alpha_{s}}{\pi}\right)^{2}\frac{\mathrm{d}^{2}\mathcal{M}^{\text{OS},(0)}}{\mathrm{d}m^{2}}\bigg{|}_{m=\overline{m}},\end{split}$ (11)
@@
Table 2: HEFT cross section in the 5-flavour scheme and for different bottom-quark masses in the 4-flavour scheme. In the last column the cross section and the scale variation are computed with the $\overline{\text{MS}}$-mass. The results are computed for LHC @ 13 TeV using the NNPDF31_nnlo_as_0118 PDF set. The central scale is chosen at $\mu_{R}=\mu_{F}=m_{H}/2$. The scale uncertainties are determined with seven-point variation. Results for the 5FS are computed with SusHi Harlander:2012pb ; Harlander:2016hcx .
@@
To validate that the cross section is free from any enhancements due to the bottom quark mass, we computed the Higgs production cross section in top-induced gluon-gluon fusion within the HEFT framework for various bottom-quark masses [^1] We note that the rescaling factor used to define HEFT does not depend on the flavour scheme in use. As indicated in eq. (2), only the effect of a finite top-quark mass is taken into account in the rescaling.. The results are presented in tab. 2. We found that the 4FS results for extremely low bottom quark masses are in very good agreement with the 5FS result, and show no sign of divergence. The findings also highlight the sizeable impact of finite-quark-mass effects, even if the quark does not directly couple to the Higgs, as the cross section is shifted by more than 2% for realistic bottom-quark masses (last two columns of tab. 2). This few percent effect seems to be in rough agreement with the order of magnitude of this effect in Higgs+jet production Pietrulewicz:2023dxt .

### v2

@@ 1 Introduction
Since the top quark is by far the heaviest quark in the SM, it also has the largest impact on the cross section, hence Higgs couplings to lighter quarks are often neglected. Moreover, the top quark’s large mass enables the use of the heavy-top limit (HTL) approximation, where the top quark is treated as infinitely massive. This method is valid up to power corrections of order $m_{H}^{2}/m_{t}^{2}$, and significantly streamlines computations by reducing the loop order by one and removing the top-quark mass as a scale. These simplifications have allowed for calculations at next-to-leading Dawson:1990zj ; Djouadi:1991tka , next-to-next-to-leading Ravindran:2002dc ; Harlander:2002wh ; Anastasiou:2002yz , and next-to-next-to-next-to-leading Anastasiou:2015vya ; Mistlberger:2018etf ; Anastasiou:2016cez order in $\alpha_{s}$ (NLO, NNLO, $\text{N}^{3}$), yielding state-of-the-art predictions with a theoretical uncertainty of approximately $6\%$. Studies have shown that the gluon-gluon-fusion channel receives particularly large perturbative corrections. In fact, the cross section is more than doubled going from LO to NLO, and NNLO corrections remain substantial, contributing roughly 20% to the total cross section. Because of the large size of the corrections, accounting for the previously omitted finite top-quark mass is strongly advisable also at higher perturbative orders. In Higgs effective field theory (HEFT), finite-top-quark-mass effects are included by rescaling the HTL cross section to recover the LO result with finite top-quark mass:
@@
Finite-quark-mass effects can also be leveraged to determine the Yukawa couplings to the respective quarks. The sensitivity of the measurements can be significantly enhanced by additionally considering differential cross sections, e.g. the Higgs-$p_{T}$ Bishara:2016jga ; Bonner:2016sdg or the Higgs-rapidity Soreq:2016rae distribution. This is particularly useful for lighter quarks like the charm quark. CMS studies CMS:2018gwt ; CMS:2023gjz have found an upper bound for the charm Yukawa coupling of about five times its SM value, by fitting NLO predictions to the measured distributions. Meanwhile, NNLO predictions for $p_{T}$ distributions with finite top and bottom mass have become available Grazzini:2013mca ; Buschmann:2014sia ; Melnikov:2016emg ; Lindert:2017pky ; Caola:2018zye ; Bonciani:2022jmb and could potentially yield improved bounds. Rapidity distributions on the other hand are even more challenging, as they require a contribution from the three-loop virtual corrections absent in the $p_{T}$ distribution. Our study aims to bridge this gap, potentially enhancing the sensitivity for future analyses of this kind.
@@ 2 Computational Details
with $\hat{s},\hat{t}$ and $\hat{u}$ the usual partonic Mandelstam variables, remains challenging. Since the computation time needed for the reduction of the amplitudes strongly scales with the polynomial degree of the coefficients in front of the master integrals, it is advantageous to search for a basis in which the coefficients become small in size. We heuristically explore different choices of bases at fixed values for the kinematic variables and perform a full reduction to our favored basis of master integrals retaining the complete symbolic dependence.
Furthermore, we use a fixed ratio of the quark and Higgs mass square, $m_{q}^{2}/m_{H}^{2}$, to eliminate one of the scales in favour of a rational number, which further simplifies the rational-function coefficients in front of the master integrals. In fact, we compute the master integrals and the complete amplitudes for five different bottom-quark masses and four different top-quark masses, listed in tab. 1. After the reduction to master integrals, we derive a set of coupled differential equations. The integrals for the boundary condition at $x\rightarrow\infty$ can be solved with the help of the large-mass expansion. We then transport the solution to physical values of $x$, by solving the differential equations numerically using the Bulirsch-Stoer algorithm (see fig. 1).
@@
This is done multiple times for many values of $z$ and a fixed value of $\lambda$ close to the symmetry axis of $\lambda=1/2$. Subsequently, we solve the differential equations in $\lambda$ to map out the entire physical plane, marked red in fig. 1. We thus obtained three-dimensional grid points of the amplitude which we use to construct cubic spline fits during the Monte Carlo integration. To assess the accuracy of the interpolation method, we removed specific grid points—in this case we removed all respective points corresponding to $\overline{m}_{b}(m_{H}),\overline{m}_{b}(m_{H}/2),\overline{m}_{b}(m_{H}/4)$ or $m_{t}$—and reran the Monte Carlo simulation. The discrepancies between the recalculated cross sections and the original ones were below the Monte Carlo uncertainties, confirming the robustness of our interpolation method. We note that we only calculated the interference of all bottom-quark amplitudes with the OS-renormalised, as well as with the $\overline{\mathrm{MS}}$-renormalised top-quark amplitudes at the same renormalisation scale, to construct grids for cubic-spline interpolation.
@@
To further improve the quality of the spline fits, we subtract all kinematic singularities arising from soft or collinear phase space regions from the amplitudes and add them back only after performing the interpolation. The subtraction of kinematic singularities can be simplified by first subtracting the amplitude square of the rescaled HTL (rHTL) defined by
$\braket{\mathcal{M}^{(0)}_{\text{rHTL}}}{\mathcal{M}^{(1)}_{\text{rHTL}}}\equiv\frac{\sigma^{\text{LO}}_{i}}{\sigma^{\text{LO}}_{\text{HTL}}}\braket{\mathcal{M}^{(0)}_{\text{HTL}}}{\mathcal{M}^{(1)}_{\text{HTL}}}.$ (4)
$\braket{\mathcal{M}^{(0)}}{\mathcal{M}^{(1)}}$ denotes the usual spin- and color-averaged interference of the LO and NLO amplitudes. $\sigma_{i}$ refers to either the cross section with finite top-quark mass, in which case the rHTL is identical to HEFT, or the top-bottom interference contribution. The regulated NLO contribution of the squared matrix element are given by
@@
$\begin{split}&\braket{\mathcal{M}^{(0)}_{gg\rightarrow gH}}{\mathcal{M}^{(1)}_{gg\rightarrow gH}}|_{\text{regulated}}\equiv\braket{\mathcal{M}^{(0)}_{gg\rightarrow gH}}{\mathcal{M}^{(1)}_{gg\rightarrow gH}}-\braket{\mathcal{M}^{(0)}_{gg\rightarrow gH,\text{rHTL}}}{\mathcal{M}^{(1)}_{gg\rightarrow gH,\text{rHTL}}}\\ &\quad+\braket{\mathcal{M}^{(0)}_{gg\rightarrow H}}{\mathcal{M}^{(1)}_{gg\rightarrow H}-\mathcal{M}^{(1)}_{gg\rightarrow H,\text{rHTL}}}\times\begin{cases}\frac{8\pi\alpha_{s}}{\hat{t}}\braket{P_{gg}^{(0)}\!\left(\frac{\hat{s}}{\hat{s}+\hat{u}}\right)}\quad&\text{if }|\hat{t}|<|\hat{u}|\\ \frac{8\pi\alpha_{s}}{\hat{u}}\braket{P_{gg}^{(0)}\!\left(\frac{\hat{s}}{\hat{s}+\hat{t}}\right)}\quad&\text{if }|\hat{u}|<|\hat{t}|\end{cases},\\[5.69046pt] &\braket{\mathcal{M}^{(0)}_{qg\rightarrow qH}}{\mathcal{M}^{(1)}_{qg\rightarrow qH}}|_{\text{regulated}}\equiv\braket{\mathcal{M}^{(0)}_{qg\rightarrow qH}}{\mathcal{M}^{(1)}_{qg\rightarrow qH}}-\braket{\mathcal{M}^{(0)}_{qg\rightarrow qH,\text{rHTL}}}{\mathcal{M}^{(1)}_{qg\rightarrow qH,\text{rHTL}}}\\ &\quad-\braket{\mathcal{M}^{(0)}_{gg\rightarrow H}}{\mathcal{M}^{(1)}_{gg\rightarrow H}-\mathcal{M}^{(1)}_{gg\rightarrow H,\text{rHTL}}}\times\frac{8\pi\alpha_{s}}{\hat{t}}\braket{P_{q\bar{q}}^{(0)}\!\left(\frac{\hat{s}}{\hat{s}+\hat{u}}\right)},\\[5.69046pt] &\braket{\mathcal{M}^{(0)}_{q\bar{q}\rightarrow gH}}{\mathcal{M}^{(1)}_{q\bar{q}\rightarrow gH}}|_{\text{regulated}}\equiv\braket{\mathcal{M}^{(0)}_{q\bar{q}\rightarrow gH}}{\mathcal{M}^{(1)}_{q\bar{q}\rightarrow gH}}-\braket{\mathcal{M}^{(0)}_{q\bar{q}\rightarrow gH,\text{rHTL}}}{\mathcal{M}^{(1)}_{q\bar{q}\rightarrow gH,\text{rHTL}}},\\ \end{split}$ (5)
where the averaged splitting functions are defined as
$\braket{P_{gg}^{(0)}\!\left(z\right)}=2C_{A}\left(\frac{z}{1-z}+\frac{1-z}{z}+z(1-z)\right)\text{ and }\braket{P_{q\bar{q}}^{(0)}(z)}=T_{F}\left(1-2z(1-z)\right).$ (6)
By subtracting the one-loop squared amplitude in the rHTL $\braket{\mathcal{M}^{(0)}_{i\rightarrow f,\text{rHTL}}}{\mathcal{M}^{(1)}_{i\rightarrow f,\text{rHTL}}}$, we already capture most of the singularities, namely all singularities arising from one-loop soft functions or collinear splitting functions. These would only act on the difference between the LO form factor and its rHTL counter part, which vanishes by design (see eq. (6)). The only singularities which are not yet removed stem from either a soft and collinear splitting of the NLO form factors. These singularities are then subtracted with the corresponding tree-level splitting function (see second and fourth line of eq. (5)). Note that the splitting function also removes the soft divergences. The regulated squared matrix element is now free of non-integrable kinematic singularities.
@@ 2.1 $\overline{\text{MS}}$ scheme
$Z_{m}^{\overline{\text{MS}}}(\mu)\overline{m}(\mu)=Z_{m}^{\text{OS}}m^{\text{OS}},$ (7)
@@
$m^{\text{OS}}=\overline{m}(\mu)\left(1+c_{1}(\mu)\frac{\alpha_{s}(\mu)}{\pi}+c_{2}(\mu)\left(\frac{\alpha_{s}(\mu)}{\pi}\right)^{2}+\mathcal{O}(\alpha^{3})\right).$ (8)
@@
$\begin{split}c_{1}(&\mu)=C_{F}\left(1-\frac{3}{4}\ln\!\left(\frac{\overline{m}(\mu)^{2}}{\mu^{2}}\right)\right),\\ c_{2}(&\mu)=-\frac{1}{384}C_{F}\bigg{(}C_{A}\left(-132\ln^{2}\!\left(\frac{\overline{m}(\mu)^{2}}{\mu^{2}}\right)+740\ln\!\left(\frac{\overline{m}(\mu)^{2}}{\mu^{2}}\right)+144\zeta(3)-1111+\pi^{2}\left(32-96\ln(2)\right)\right)\\ &\hskip 28.45274pt+3C_{F}\bigg{(}-36\ln^{2}\!\left(\frac{\overline{m}(\mu)^{2}}{\mu^{2}}\right)-36\ln\!\left(\frac{\overline{m}(\mu)^{2}}{\mu^{2}}\right)-96\zeta(3)+71+8\pi^{2}(8\ln(2)-5)\bigg{)}\\ &\hskip 28.45274pt+4n_{l}T_{F}\left(12\ln^{2}\!\left(\frac{\overline{m}(\mu)^{2}}{\mu^{2}}\right)-52\ln\!\left(\frac{\overline{m}(\mu)^{2}}{\mu^{2}}\right)+8\pi^{2}+71\right)\bigg{)}\\ &-\sum_{i}\frac{1}{96}C_{F}T_{F}\bigg{(}12\ln^{2}\!\left(\frac{\overline{m}(\mu)^{2}}{\mu^{2}}\right)-24\ln\!\left(\frac{\overline{m}(\mu)^{2}}{\mu^{2}}\right)\ln\!\left(\frac{\overline{m}_{i}(\mu)^{2}}{\mu^{2}}\right)-52\ln\!\left(\frac{\overline{m}(\mu)^{2}}{\mu^{2}}\right)+32\ln\!\left(\frac{\overline{m}_{i}(\mu)^{2}}{\mu^{2}}\right)\\ &\hskip 28.45274pt+48(x_{i}-1)^{2}\left(x_{i}^{2}+x_{i}+1\right)\text{Li}_{2}(1-x_{i})-48\left(x_{i}^{4}+x_{i}^{3}+x_{i}+1\right)\left(\text{Li}_{2}(-x_{i})+\ln\!(x_{i})\ln\!(x_{i}+1)\right)\\ &\hskip 28.45274pt+48x_{i}^{4}\ln^{2}(x_{i})+48x_{i}^{2}\ln(x_{i})-16\pi^{2}(x_{i}^{3}+x_{i})+72x_{i}^{2}+71\bigg{)},\end{split}$ (10)
where we defined $x_{i}=\frac{\overline{m}_{i}(\mu)}{\overline{m}(\mu)}$. The sum runs over all heavy-quark flavours which are decoupled from the running of $\alpha_{s}$. This includes $\overline{m}$ itself in case the corresponding quark is decoupled. Subsequently, the amplitudes can simply be expanded to find the relation between the two schemes:
$\begin{split}&\mathcal{M}^{\overline{\text{MS}}}(\overline{m}(\mu))=\mathcal{M}^{\text{OS}}(\overline{m}(\mu))+\delta\mathcal{M}(\overline{m}(\mu)),\\ &\delta\mathcal{M}^{(1)}(\overline{m}(\mu))=\overline{m}(\mu)c_{1}(\mu)\frac{\alpha_{s}(\mu)}{\pi}\frac{\mathrm{d}\mathcal{M}^{\text{OS},(0)}}{\mathrm{d}m}\bigg{|}_{m=\overline{m}(\mu)},\\ &\delta\mathcal{M}^{(2)}(\overline{m}(\mu))=\overline{m}(\mu)\left[c_{1}(\mu)\frac{\alpha_{s}(\mu)}{\pi}\frac{\mathrm{d}\mathcal{M}^{\text{OS},(1)}}{\mathrm{d}m}\bigg{|}_{m=\overline{m}(\mu)}+c_{2}(\mu)\left(\frac{\alpha_{s}(\mu)}{\pi}\right)^{2}\frac{\mathrm{d}\mathcal{M}^{\text{OS},(0)}}{\mathrm{d}m}\bigg{|}_{m=\overline{m}(\mu)}\right]\\ &\hskip 227.62204pt+\frac{1}{2}\left(\overline{m}(\mu)c_{1}(\mu)\frac{\alpha_{s}(\mu)}{\pi}\right)^{2}\frac{\mathrm{d}^{2}\mathcal{M}^{\text{OS},(0)}}{\mathrm{d}m^{2}}\bigg{|}_{m=\overline{m}(\mu)},\end{split}$ (11)
@@
Table 2: HEFT cross section in the 5-flavour scheme and for different bottom-quark masses in the 4-flavour scheme. In the last column the cross section and the scale variation are computed with the $\overline{\text{MS}}$-mass. The results are computed for LHC @ 13 TeV using the NNPDF31_nnlo_as_0118 PDF set in the 5FS and the NNPDF31_nnlo_as_0118_nf_4 PDF set in the 4FS. The central scale is chosen at $\mu_{R}=\mu_{F}=m_{H}/2$. The scale uncertainties are determined with seven-point variation. Results for the 5FS are computed with SusHi Harlander:2012pb ; Harlander:2016hcx .
@@
To validate that the cross section is free from any enhancements due to the bottom quark mass, we computed the Higgs production cross section in top-induced gluon-gluon fusion within the HEFT framework for various bottom-quark masses [^1] We note that the rescaling factor used to define HEFT does not depend on the flavour scheme in use. As indicated in eq. (2), only the effect of a finite top-quark mass is taken into account in the rescaling.. The results are presented in tab. 2. At LO, the only difference between the 4FS and 5FS is the used PDF set which is why there is only a marginal discrepancy between the cross sections. At NLO, the bottom quark does not appear in any 4FS amplitudes yet, hence the cross section is still independent of the bottom quark mass. In the 5FS however, the $bg\longrightarrow Hb$ channel is now opening up, causing a more significant deviation from the 4FS than at LO. Finally at NNLO the 4FS cross sections become dependent on the bottom quark mass through the diagrams depicted in fig. 2, 3 and 4 with the top quark, represented by the thick fermion lines, integrated out, i.e. all top-quark loops shrink down to effective vertices. We found that the 4FS results for extremely low bottom-quark masses are numerically very similar to the 5FS result, and we see no sign of divergence. The findings also highlight the sizeable impact of finite-quark-mass effects, even if the quark does not directly couple to the Higgs, as the cross section is shifted by more than 2% for realistic bottom-quark masses (last two columns of tab. 2). This few percent effect seems to be in rough agreement with the order of magnitude of this effect in Higgs+jet production Pietrulewicz:2023dxt .